# Design of efficient high-order immersed metagratings using an evolutionary algorithm

### v2

@@ 1 Introduction
High-resolution spectrometers are essential in ground-based and space-based astronomy, where compact and high-performance optical components are critical. Immersed diffraction gratings are particularly favourable due to their ability to provide high resolving power while maintaining a compact volume. They are currently used in instruments such as SPIFFIER on the VLT [george2016making], and are planned for first-light instruments on upcoming extremely large telescopes, such as METIS for the Extremely Large Telescope (ELT) [agocs2020ge] and GMTNIRS for the Giant Magellan Telescope (GMT) [jaffe2016gmtnirs], as well as in space-based missions like Sentinel-5 [kohlhaas2019manufacturing].
@@
The power concentrated into the diffraction angle ($\theta_{d}$) depends on the groove shape and the incidence angle ($\theta_{i}$) [soskind2011field]. Specifically, the blaze angle $\theta_{b}$ of the grating (see Figure 1 (a) and (b)) is related to the diffraction and incidence angles by the blazing condition. This is given by $\theta_{b}=\theta_{i}\pm\theta_{d}$. A typical strategy to obtain a blazed grating profile is through wet etching into a crystalline material, where the resulting etching angle (formed by the two groove surfaces, $\eta$ in Figure 1 (b)) is fixed by the crystal axes of the material [philipsen2005anisotropy]. For example, in KOH (potassium hydroxide) etching of amorphous silicon (a-Si), the etching angle $\eta$ is fixed to $\sim 70.5^{\circ}$ [kohlhaas2019manufacturing]. This limitation affects the efficiency of the grating since an idealized groove shape is not realized. This can be mitigated using a more precise machining process, such as single-point diamond turning (SPDT) machining as used for METIS’s mid-infrared Ge grating [agocs2020ge]. However, SPDT struggles with shorter periods (< 5-10 $\mu\mathrm{m}$), which is required for optical, near-infrared and infrared applications.
Metasurfaces are used for robust wavefront control [yu2014flat] and have been applied in areas such as beam steering [berini2022opticalBS], waveplates [ding2015broadbandWP], and diffraction gratings operating in both transmission [erdmann2017developmentGr] and reflection [peng2011reflective]. Blazed diffraction metagratings based on binary surfaces, designed to mimic the blazed $2\pi$ phase profile, have also been demonstrated in optical [kitt2015visibleGr, lalanne1998blazed], infrared [lee2004blazed], and even broadband [lee2023wide] applications. Beyond conventional diffraction control, metagratings have also been used to generate structured beams such as vortex beams [zhang2021generating], and for polarisation state generation [rubin2018polarization].
@@
Metasurfaces are implemented using lithography techniques in the form of nanopillars [lee2004blazed], ridges [lalanne1998blazed] [zeitner2023high], or both [zeitner2012high]. By varying dimensions (see Figure 1 (d)) the subwavelength structure’s ability to behave as an effective homogeneous medium [cheben2018subwavelength] can be leveraged. This is referred to as effective index dispersion engineering based on Rytov’s effective medium theory (EMT) [rytov1956electromagnetic]. A blazed profile, even when ideally implemented using metasurfaces, offers only a single degree of freedom (the blaze angle that is mimicked by the metasurfaces). This limits the ability to design higher-order gratings with minimal polarization dependence, even over narrowband spectral ranges. Additionally, as evident from the grating equation, dispersion in blazed gratings is inherently chromatic, and the blaze angle is optimized to concentrate power into a specific diffraction order at an operating wavelength. This leads to wavelength and polarization-dependent variations in diffraction efficiency. Therefore, rather than using metasurfaces strictly to emulate an idealized sawtooth blazed profile (“Blazed profile” and “sawtooth profile” are used interchangeably in this paper) for one wavelength, they may be employed more flexibly to optimize the grating’s performance directly over a wavelength range. Additionally, the modulation by EMT depends on the incident field orientation and is therefore polarization sensitive [hemmati2020applicability]. Hence, as with conventional blazed gratings [baldry2004volume], metagratings also exhibit polarization dependence. By directly optimizing the performance of the metagrating, the sensitivity to polarization and chromaticity can be minimized compared to a conventional sawtooth grating, as shown in [ans2024topology] for a $\pm 1$st order grating case.
@@
In this study, an evolutionary algorithm (EA) described in [de2021tuning] is used to optimize the design of a higher-order immersed metagrating operating in the shortwave infrared (SWIR-3) band ($2.304\,\mu\text{m}-2.405\,\mu\text{m}$). The reference grating for the Sentinel-5 mission was manufactured with KOH wet etching, which leads to smooth side walls, but constrains the groove angle [10.1117/12.461759]. Similar to the Sentinel-5 grating, the metagrating is immersed in silicon ($Si$) and coated with aluminum ($Al$) to operate in reflection, but uses nanopillars for the grating operation. Figure 1 compares schematics of such a grating with a conventional blazed sawtooth grating. Unlike the conventional approach that mimics a single blaze angle, the metagrating is directly optimized over a wavelength range to maximize diffraction efficiency and minimize polarization sensitivity. The motivation behind using an optimization algorithm and the principles behind the design are explained in Section 2.1. The design problem, the optimization algorithm, the figure of merit and the simulation method are described in Section 2.2. Results from the optimization process and the optimal metagrating design are presented in subsection 3.1. The performance of the optimized design—including its diffraction efficiency, polarization dependence, and dispersion characteristics—is shown in Section 3.2. The limitations of the current optimization approach and the manufacturing considerations are also discussed in Sections 3.1 and 3.2, respectively. Finally, the limitations of this study and directions for future work with a summary are outlined in Section 4.
@@ 2.1 Design principle
The design is constrained to have a fixed period $P$ (see Figure 1) of $2.07\,\mu\text{m}$, with an incidence angle $\theta_{i}=62.6^{\circ}$, diffraction order m=5m=-5, and consequently a diffraction angle $\theta_{d}=-49.8^{\circ}$. These constraints are based on the design parameters of the Sentinel-5 SWIR-3 (centered at $\lambda_{\text{mean}}=2.345\,\mu\text{m}$ with $\sim 4\%$ bandwidth) blazed grating described in [kohlhaas2019manufacturing]. This establishes a baseline for comparing the proposed immersed metagrating to a conventional immersed blazed grating.
The metagrating, in its simplest form, consists of a periodically arranged subwavelength structure. For a structure with period $\Lambda$ composed of alternating materials with complex refractive indices $\tilde{n}_{1}$ and $\tilde{n}_{2}$, and respective widths $a$ and $b$, operating at a free-space wavelength $\lambda$, the effective refractive index $n_{\text{eff}}$ is governed by the full formalism of Rytov’s theory. The expressions for the two orthogonal polarizations are given by Equations 1 and 2, respectively [rytov1956electromagnetic, hemmati2020applicability].
@@
Because the transcendental EMT equations (Equations 12) involve tangent functions with periodic singularities, they admit an infinite family of solutions for $n_{\text{TE}}^{\rm EMT}$ and $n_{\text{TM}}^{\rm EMT}$ that depend on the materials ($n_{1},\,n_{2}$), wavelength($\lambda$) and the design parameters($\Lambda,\,a,\,b$). In practice, only the principal (first-order) root can be retained [hemmati2020applicability, cheben2018subwavelength], but is only valid when the grating period $\Lambda$ is vanishingly small compared to the free-space wavelength $\lambda$. In our design, the period $2.07\,\mu\mathrm{m}$ is divided into five sub-periods of $0.414\,\mu\mathrm{m}$ to meet fabrication aspect-ratio constraints; although $0.414\,\mu\mathrm{m}<2.345\,\mu\mathrm{m}$, the ratio $(\Lambda/\lambda\approx 0.18$) is large enough for the first order-approximation to hold. Additionally, the nonzero extinction coefficients of silicon and aluminum would shift the roots. Consequently, there is no closed-form guarantee that the simple first-order EMT solution will maximize the grating efficiency (defined here as diffraction efficiency $\times$ reflection efficiency). Moreover, to use the Rytov-based EMT to approximate a profile, an ideal profile is required that performs well across the spectral range and accounts for polarization effects. A conventional blazed (sawtooth) profile is typically optimized at a single wavelength within the spectral band, and its performance is therefore not uniform across the band. Furthermore, the blazed profile does not account for polarization. To address these limitations and to exploit subwavelength structuring to engineer an arbitrary effective medium, an evolutionary algorithm (EA) is employed. The EA effectively optimizes the geometry to maximize the overall grating efficiency while also minimizing sensitivity to polarization and wavelength variations through an appropriately defined figure of merit (FOM). The underlying principle enabling this approach is that the optimizer discovers metasurfaces with a refractive index profile that is well adapted to the problem at hand.
@@ 2.2 Simulation and optimization of design
As for the design variables, each $P\times P$ unit cell of the grating is represented by a $5\times 5$ grid of subcells. As mentioned previously, the $5\times 5$ grid of subcells comes from the period divided into 5 parts, so as to balance the subwavelength regime and manufacturing tolerance. In the xx–direction, these subcells vary in width and length (varying fill-factor) to approximate the target phase profile; in the zz–direction, the pattern is kept uniform (constant fill-factor), yielding an array of nanopillars (refer to figure 2). This effectively ensures that the diffraction occurs only along one direction and only the fundamental order is excited in the other direction. Furthermore, the metasurfaces are limited to nanopillars rather than ridges. It has been shown before that using nanopillars rather than continuous ridges could improve the stability of the structure while keeping a low fill-factor [zeitner2010high]. All pillars share a common height $h$. The common height means the manufacturing is simplified, and the phase difference between each metasurface is purely from the fill-factor difference. Thus, the structure is parameterized by eleven degrees of freedom: the width ($b_{x_{i}}$) and length ($b_{z_{i}}$) of the 5 subcells (i = 1 - 5) along the $x$ direction, and the pillars’ height ($h$) (again refer to figure 2). This yields the design vector $\mathbf{x}=[b_{x_{1}},b_{z_{1}},\dots,b_{x_{5}},b_{z_{5}},H]$ of length $n_{\mathbf{x}}=11$. A numeric optimizer $\mathbf{x}^{*}$ is found which maximizes our figure of merit (FOM) $f:\mathbf{S}\to\mathbb{R}$:
@@
Since obtaining a closed-form analytical solution is infeasible, it is considered a black-box optimization problem [audet2017derivative]. As such, a heuristic optimization algorithm can be applied, which utilizes a stochastic process to optimize $\mathbf{x}$ by considering only the value of $f(\mathbf{x})$. Here, the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [hansen2001completely] is used. The CMA-ES is an EA designed for solving challenging continuous black-box optimization problems and is considered state-of-the-art in heuristic optimization [hansen2010comparing]. The algorithm can be considered a second-order method, which estimates the parameters of a multivariate Gaussian distribution $\mathcal{N}(\mathbf{m},\sigma^{2}\mathbf{C})$ that it uses to sample new $\mathbf{x}$. At every iteration, the algorithm samples $k$ candidate solutions of which $l<k$ are selected (Note that in standard CMA-ES literature $\mu$ and $\lambda$ are used for $l$ and $k$ resp.). The method then updates the parameters of the distribution, such that the likelihood of sampling the $l$ selected steps again is maximized. In essence, this creates a sampling distribution that forms a hyperellipsoid pointed towards the direction of most progress.
@@
Over the years, several extensions to the CMA-ES have been proposed to enhance its search performance, many of which have been implemented within the Modular CMA-ES framework [denobel2021tuning], which is utilized in this work. Specifically, the variant used here includes active covariance matrix updates [jastrebski2006improving] and mirrored sampling [auger2011mirrored]. Since CMA-ES is defined as a minimization algorithm, the function $f(\mathbf{x})$ is optimized internally. In this study, CMA-ES is applied as a practical optimization tool rather than investigated as an algorithmic contribution. It was chosen for its robustness in non-convex black-box problems, comparable to the present design task [smaoui2018combination, ghahremani2020fast, hamdani2019optimization]. Other stochastic methods (e.g., PSO, DE, Bayesian optimization) could also be used, but a full benchmarking is beyond the scope of this work. CMA-ES serves here as a representative, well-established optimizer with demonstrated reliability [hansen2010comparing]. The implementation follows the Modular CMA-ES framework (See: [denobel2021tuning]) , with population size $k=11$, parent number $l=5$, initial step size $\sigma_{0}=0.6$, and a maximum of $N_{\text{eval}}=500$ evaluations. The initial mean $\mathbf{m}_{0}$ was placed at the center of the design space with isotropic covariance $\mathbf{C}_{0}=\mathbf{I}$. The algorithm employs the standard (non-elitist) selection scheme with weighted recombination, adhering to the canonical CMA-ES [hansen2001completely]. The total computational cost is dominated by the FDTD simulations, each of which takes $\pm$ 12 minutes per evaluation, while the optimizer’s overhead is negligible. Because CMA-ES is a stochastic, derivative-free method, conventional convergence-rate metrics are not applicable; convergence is instead indicated by the stabilization of the objective value.
@@ 3.1 CMA-ES optimization results
All the pillars have feature sizes just above $150\,\mathrm{nm}$, which is well within the current e-beam lithography capabilities for Si, as discussed in [zeitner2023potential]. The designed metagrating can even be prototyped with deep UV photo-lithography (DUV-PL) with methods to improve for smaller feature sizes ($\sim 100\,\mathrm{nm}$) at reduced precision as demonstrated in [hu2020cmos]. The largest aspect ratio is $\sim 1.3$, which further relaxes the manufacturing constraints. The Al filling in the etched pillars and coating could be done by sputtering.

### v3

@@ 1 Introduction
High-resolution spectrometers are essential in ground-based and space-based astronomy, where compact and high-performance optical components are critical. Immersed diffraction gratings are particularly favourable due to their ability to provide high resolving power while maintaining a compact volume. They are currently used in instruments such as SPIFFIER on the VLT [12], and are planned for first-light instruments on upcoming extremely large telescopes, such as METIS for the Extremely Large Telescope (ELT) [1] and GMTNIRS for the Giant Magellan Telescope (GMT) [20], as well as in space-based missions like Sentinel-5 [25].
@@
The power concentrated into the diffraction angle ($\theta_{d}$) depends on the groove shape and the incidence angle ($\theta_{i}$) [37]. Specifically, the blaze angle $\theta_{b}$ of the grating (see Figure 1 (a) and (b)) is related to the diffraction and incidence angles by the blazing condition. This is given by $\theta_{b}=\theta_{i}\pm\theta_{d}$. A typical strategy to obtain a blazed grating profile is through wet etching into a crystalline material, where the resulting etching angle (formed by the two groove surfaces, $\eta$ in Figure 1 (b)) is fixed by the crystal axes of the material [32]. For example, in KOH (potassium hydroxide) etching of amorphous silicon (a-Si), the etching angle $\eta$ is fixed to $\sim 70.5^{\circ}$ [25]. This limitation affects the efficiency of the grating since an idealized groove shape is not realized. This can be mitigated using a more precise machining process, such as single-point diamond turning (SPDT) machining as used for METIS’s mid-infrared Ge grating [1]. However, SPDT struggles with shorter periods (< 5-10 $\mu\mathrm{m}$), which is required for optical, near-infrared and infrared applications.
Metasurfaces are used for robust wavefront control [40] and have been applied in areas such as beam steering [6], waveplates [10], and diffraction gratings operating in both transmission [11] and reflection [31]. Blazed diffraction metagratings based on binary surfaces, designed to mimic the blazed $2\pi$ phase profile, have also been demonstrated in optical [24, 26], infrared [28], and even broadband [27] applications. Beyond conventional diffraction control, metagratings have also been used to generate structured beams such as vortex beams [45], and for polarisation state generation [33].
@@
Metasurfaces are implemented using lithography techniques in the form of nanopillars [28], ridges [26] [43], or both [44]. By varying dimensions (see Figure 1 (d)) the subwavelength structure’s ability to behave as an effective homogeneous medium [7] can be leveraged. This is referred to as effective index dispersion engineering based on Rytov’s effective medium theory (EMT) [34]. A blazed profile, even when ideally implemented using metasurfaces, offers only a single degree of freedom (the blaze angle that is mimicked by the metasurfaces). This limits the ability to design higher-order gratings with minimal polarization dependence, even over narrowband spectral ranges. Additionally, as evident from the grating equation, dispersion in blazed gratings is inherently chromatic, and the blaze angle is optimized to concentrate power into a specific diffraction order at an operating wavelength. This leads to wavelength and polarization-dependent variations in diffraction efficiency. Therefore, rather than using metasurfaces strictly to emulate an idealized sawtooth blazed profile (“Blazed profile” and “sawtooth profile” are used interchangeably in this paper) for one wavelength, they may be employed more flexibly to optimize the grating’s performance directly over a wavelength range. Additionally, the modulation by EMT depends on the incident field orientation and is therefore polarization sensitive [18]. Hence, as with conventional blazed gratings [5], metagratings also exhibit polarization dependence. By directly optimizing the performance of the metagrating, the sensitivity to polarization and chromaticity can be minimized compared to a conventional sawtooth grating, as shown in [2] for a $\pm 1$st order grating case.
@@
In this study, an evolutionary algorithm (EA) described in [8] is used to optimize the design of a higher-order immersed metagrating operating in the shortwave infrared (SWIR-3) band ($2.304\,\mu\text{m}-2.405\,\mu\text{m}$). The reference grating for the Sentinel-5 mission was manufactured with KOH wet etching, which leads to smooth side walls, but constrains the groove angle [29]. Similar to the Sentinel-5 grating, the metagrating is immersed in silicon ($Si$) and coated with aluminum ($Al$) to operate in reflection, but uses nanopillars for the grating operation. Figure 1 compares schematics of such a grating with a conventional blazed sawtooth grating. Unlike the conventional approach that mimics a single blaze angle, the metagrating is directly optimized over a wavelength range to maximize diffraction efficiency and minimize polarization sensitivity. The motivation behind using an optimization algorithm and the principles behind the design are explained in Section 2.1. The design problem, the optimization algorithm, the figure of merit and the simulation method are described in Section 2.2. Results from the optimization process and the optimal metagrating design are presented in subsection 3.1. The performance of the optimized design—including its diffraction efficiency, polarization dependence, and dispersion characteristics—is shown in Section 3.2. The limitations of the current optimization approach and the manufacturing considerations are also discussed in Sections 3.1 and 3.2, respectively. Finally, the limitations of this study and directions for future work with a summary are outlined in Section 4.
@@ 2.1 Design principle
The design is constrained to have a fixed period $P$ (see Figure 1) of $2.07\,\mu\text{m}$, with an incidence angle $\theta_{i}=62.6^{\circ}$, diffraction order m=5m=-5, and consequently a diffraction angle $\theta_{d}=-49.8^{\circ}$. These constraints are based on the design parameters of the Sentinel-5 SWIR-3 (centered at $\lambda_{\text{mean}}=2.345\,\mu\text{m}$ with $\sim 4\%$ bandwidth) blazed grating described in [25]. This establishes a baseline for comparing the proposed immersed metagrating to a conventional immersed blazed grating.
The metagrating, in its simplest form, consists of a periodically arranged subwavelength structure. For a structure with period $\Lambda$ composed of alternating materials with complex refractive indices $\tilde{n}_{1}$ and $\tilde{n}_{2}$, and respective widths $a$ and $b$, operating at a free-space wavelength $\lambda$, the effective refractive index $n_{\text{eff}}$ is governed by the full formalism of Rytov’s theory. The expressions for the two orthogonal polarizations are given by Equations 1 and 2, respectively [34, 18].
@@
Because the transcendental EMT equations (Equations 12) involve tangent functions with periodic singularities, they admit an infinite family of solutions for $n_{\text{TE}}^{\rm EMT}$ and $n_{\text{TM}}^{\rm EMT}$ that depend on the materials ($n_{1},\,n_{2}$), wavelength($\lambda$) and the design parameters($\Lambda,\,a,\,b$). In practice, only the principal (first-order) root can be retained [18, 7], but is only valid when the grating period $\Lambda$ is vanishingly small compared to the free-space wavelength $\lambda$. In our design, the period $2.07\,\mu\mathrm{m}$ is divided into five sub-periods of $0.414\,\mu\mathrm{m}$ to meet fabrication aspect-ratio constraints; although $0.414\,\mu\mathrm{m}<2.345\,\mu\mathrm{m}$, the ratio $(\Lambda/\lambda\approx 0.18$) is large enough for the first order-approximation to hold. Additionally, the nonzero extinction coefficients of silicon and aluminum would shift the roots. Consequently, there is no closed-form guarantee that the simple first-order EMT solution will maximize the grating efficiency (defined here as diffraction efficiency $\times$ reflection efficiency). Moreover, to use the Rytov-based EMT to approximate a profile, an ideal profile is required that performs well across the spectral range and accounts for polarization effects. A conventional blazed (sawtooth) profile is typically optimized at a single wavelength within the spectral band, and its performance is therefore not uniform across the band. Furthermore, the blazed profile does not account for polarization. To address these limitations and to exploit subwavelength structuring to engineer an arbitrary effective medium, an evolutionary algorithm (EA) is employed. The EA effectively optimizes the geometry to maximize the overall grating efficiency while also minimizing sensitivity to polarization and wavelength variations through an appropriately defined figure of merit (FOM). The underlying principle enabling this approach is that the optimizer discovers metasurfaces with a refractive index profile that is well adapted to the problem at hand.
@@ 2.2 Simulation and optimization of design
As for the design variables, each $P\times P$ unit cell of the grating is represented by a $5\times 5$ grid of subcells. As mentioned previously, the $5\times 5$ grid of subcells comes from the period divided into 5 parts, so as to balance the subwavelength regime and manufacturing tolerance. In the xx–direction, these subcells vary in width and length (varying fill-factor) to approximate the target phase profile; in the zz–direction, the pattern is kept uniform (constant fill-factor), yielding an array of nanopillars (refer to figure 2). This effectively ensures that the diffraction occurs only along one direction and only the fundamental order is excited in the other direction. Furthermore, the metasurfaces are limited to nanopillars rather than ridges. It has been shown before that using nanopillars rather than continuous ridges could improve the stability of the structure while keeping a low fill-factor [41]. All pillars share a common height $h$. The common height means the manufacturing is simplified, and the phase difference between each metasurface is purely from the fill-factor difference. Thus, the structure is parameterized by eleven degrees of freedom: the width ($b_{x_{i}}$) and length ($b_{z_{i}}$) of the 5 subcells (i = 1 - 5) along the $x$ direction, and the pillars’ height ($h$) (again refer to figure 2). This yields the design vector $\mathbf{x}=[b_{x_{1}},b_{z_{1}},\dots,b_{x_{5}},b_{z_{5}},H]$ of length $n_{\mathbf{x}}=11$. A numeric optimizer $\mathbf{x}^{*}$ is found which maximizes our figure of merit (FOM) $f:\mathbf{S}\to\mathbb{R}$:
@@
Since obtaining a closed-form analytical solution is infeasible, it is considered a black-box optimization problem [3]. As such, a heuristic optimization algorithm can be applied, which utilizes a stochastic process to optimize $\mathbf{x}$ by considering only the value of $f(\mathbf{x})$. Here, the Covariance Matrix Adaptation Evolution Strategy (CMA-ES) [17] is used. The CMA-ES is an EA designed for solving challenging continuous black-box optimization problems and is considered state-of-the-art in heuristic optimization [16]. The algorithm can be considered a second-order method, which estimates the parameters of a multivariate Gaussian distribution $\mathcal{N}(\mathbf{m},\sigma^{2}\mathbf{C})$ that it uses to sample new $\mathbf{x}$. At every iteration, the algorithm samples $k$ candidate solutions of which $l<k$ are selected (Note that in standard CMA-ES literature $\mu$ and $\lambda$ are used for $l$ and $k$ resp.). The method then updates the parameters of the distribution, such that the likelihood of sampling the $l$ selected steps again is maximized. In essence, this creates a sampling distribution that forms a hyperellipsoid pointed towards the direction of most progress.
@@
Over the years, several extensions to the CMA-ES have been proposed to enhance its search performance, many of which have been implemented within the Modular CMA-ES framework [9], which is utilized in this work. Specifically, the variant used here includes active covariance matrix updates [21] and mirrored sampling [4]. Since CMA-ES is defined as a minimization algorithm, the function $f(\mathbf{x})$ is optimized internally. In this study, CMA-ES is applied as a practical optimization tool rather than investigated as an algorithmic contribution. It was chosen for its robustness in non-convex black-box problems, comparable to the present design task [35, 13, 15]. Other stochastic methods (e.g., PSO, DE, Bayesian optimization) could also be used, but a full benchmarking is beyond the scope of this work. CMA-ES serves here as a representative, well-established optimizer with demonstrated reliability [16]. The implementation follows the Modular CMA-ES framework (See: [9]) , with population size $k=11$, parent number $l=5$, initial step size $\sigma_{0}=0.6$, and a maximum of $N_{\text{eval}}=500$ evaluations. The initial mean $\mathbf{m}_{0}$ was placed at the center of the design space with isotropic covariance $\mathbf{C}_{0}=\mathbf{I}$. The algorithm employs the standard (non-elitist) selection scheme with weighted recombination, adhering to the canonical CMA-ES [17]. The total computational cost is dominated by the FDTD simulations, each of which takes $\pm$ 12 minutes per evaluation, while the optimizer’s overhead is negligible. Because CMA-ES is a stochastic, derivative-free method, conventional convergence-rate metrics are not applicable; convergence is instead indicated by the stabilization of the objective value.
@@ 3.1 CMA-ES optimization results
All the pillars have feature sizes just above $150\,\mathrm{nm}$, which is well within the current e-beam lithography capabilities for Si, as discussed in [42]. The designed metagrating can even be prototyped with deep UV photo-lithography (DUV-PL) with methods to improve for smaller feature sizes ($\sim 100\,\mathrm{nm}$) at reduced precision as demonstrated in [19]. The largest aspect ratio is $\sim 1.3$, which further relaxes the manufacturing constraints. The Al filling in the etched pillars and coating could be done by sputtering.